please someone help me to simplify?
9
over
18
over
3
over
15

Answer by checkley75(3666)

(9/18)/(3/15)
9/18*15/3
OR
9*15/18*3 NOW REDUCE
9/3*15/18
3*5/6
15/6
2.5 ANSWER
---------------------------------
AN ALTERNTE APPROACH IS TO REDUCE THESE FRACTIONS TO DECIMALS & THEN MULTIPLY:
9/18=.5
15/3=5
.5*5=2.5